# Measuring Agents in Production

### v1

@@ 6.2 Evaluation Methods
We ask participants which evaluation strategies they employ, allowing multiple selections. These methods apply to both offline evaluation during development and online evaluation in production environments. Four methods dominate responses: human-in-the-loop evaluation, model-based evaluation, rule-based evaluation (heuristics or syntactic checks), and cross-referencing evaluation (verification against knowledge bases or reference datasets). The exact question and method descriptions appear in Appendix E.
Human-in-the-loop verification. The majority (74.2%) rely on manual, human-in-the-loop evaluation (Figure 10(b). These evaluations typically involve domain experts, operators, or end-users directly inspecting, testing, or validating system outputs to ensure correctness, safety, and reliability.
Human experts play a critical role during development for offline evaluations. Agent developers work directly with domain experts or target users to validate system responses. For example, medical professionals are directly involved when testing and evaluating the correctness of healthcare agent systems. In one case, a company even forgoes automated evaluation methods entirely, relying instead on human experts to provide feedback to hand-tune agent configurations for each client’s deployment environment.
Human experts also serve as verifiers during agent execution for online evaluation. Teams commonly have human experts perform final actions based on agent output, serving as a layer of guardrails. For example, one site reliability engineering agent suggests actions based on analysis across system stacks, but human experts make the final decision on what to execute. The agent does not directly modify the production environment.
Model-based evaluation. Model-based evaluation methods such as LLM-as-a-judge are the second most common approach, used by 51.6% of respondents (Figure 10(b)). While this evaluation method applies for both offline development time and online runtime evaluation, 7 of 20 interview case studies use LLM judges for online evaluation during agent execution.
Model-based evaluation does not eliminate human involvement. Figure 10(c) shows the co-occurrence of different evaluation strategies, revealing that among the 51.6% of survey respondents who use model-based evaluation, a substantial portion (38.7%) also employ Human-in-the-Loop verification. In detailed case studies, all interviewed teams using LLM-as-a-judge combine it with human review. Specifically, these teams use LLM judges to evaluate confidence in every final response, combined with human subsampling. An LLM judge continuously assesses each agent’s final response. If the judge scores above a preset confidence threshold, the output is accepted automatically. Otherwise, the request routes to human experts. Additionally, human experts sample a preset percentage (e.g., 5%) of production runs even when the LLM judge expresses high confidence, verifying correctness to ensure consistent alignment at runtime. Development teams update prompt instructions and agent configurations based on this production feedback.
Other methods. Rule-based evaluation methods and cross-referencing strategies show comparable adoption rates (41.9% and 38.7% respectively). Rule-based evaluation consists of simple logic checks such as grammar and syntax verification or domain-specific rules. For example, coding agents verify outputs through compilation checks and test suites, while analytical agents may apply domain-specific rubrics to assess output quality. Cross-referencing evaluation uses external sources for grounding and fact-checking to verify the accuracy and quality of generated answers or solutions. This includes retrieving supporting evidence from trusted knowledge bases or comparing outputs against reference datasets.
Evaluation method patterns. Co-occurrence analysis reveals that human-in-the-loop evaluation is the most common method used together with other evaluation strategies (Figure 10(c)). Practitioners anchor automated, rule-based, and cross-referencing methods around human judgment rather than relying on them in isolation.
Notably, human-in-the-loop (74.2%) and LLM-as-a-judge (51.6%) dominate compared to rule-based verification (42.9%). Based on our analysis, this pattern may suggests that production agents already handle complex tasks beyond classification, entity resolution, or pattern matching. These agents operate in domains requiring nuanced judgment where rule-based methods prove insufficient. For example, customer support voice assistance and human resource operations demand expertise beyond pattern matching, explaining why practitioners require human and LLM verification to assess output correctness.
7 4 Results: What Are The Top Challenges In Building Production Agents?
AI agents are deployed at scale, yet significant challenges remain. We investigate the friction points practitioners face most when building production agents. Our survey and detailed case studies reveal that reliability remains the primary bottleneck. Survey respondents at all agent stages rank “Core Technical Focus” as their top challenge (37.9%), far outweighing governance (3.4%) or compliance (17.2%) (see Appendix B.3). Core technical focus encompasses reliability, robustness, and scalability. This prioritization signals that practitioners currently focus on making agents work consistently and correctly.
We detail the three specific dimensions of these challenges: evaluation (Section 7.1), latency (Section 7.2), and security (Section 7.3). This gap highlights research opportunities to advance agents from bounded use cases toward broader applications.
@@ 7.1 Evaluation Challenges
Difficulties creating benchmarks. As defined in Section 6.1, benchmarks are curated sets of tasks or questions with known correct answers used for offline evaluation during development. Among interviewed teams, 5 build custom benchmarks. However, benchmark creation proves challenging for three reasons based on our case study interviews.
First, specific domains lack accessible public data. In regulated fields like insurance underwriting, the absence of public data forces teams to handcraft benchmark datasets from scratch through collaboration with domain experts and target users. Second, creating high-quality benchmarks at scale is resource-intensive and time-consuming. One interviewed team reported spending months creating an initial set of approximately 40 unique environment-problem-solution scenarios, followed by another six months to scale the dataset to roughly 100 examples. Third, benchmark creation is nearly infeasible for agent applications focused on highly customized client integrations. One interviewed voice agent team illustrates this challenge: while core features are straightforward to test without extensive data, the primary engineering effort involves client-specific customizations such as proprietary toolsets, product offerings, and localized dialogue tuning. Creating standardized benchmarks for every possible client configuration is impractical. Given these challenges, 75% of interviewed teams forgo benchmark creation entirely, instead relying on A/B testing or direct client collaboration to iterate based on human feedback until expectations are met.
Agent behavior breaks traditional software testing. Three case study teams report attempting but struggling to integrate agents into existing CI/CD pipelines. The challenge stems from agent nondeterminism and the difficulty of judging outputs programmatically. Despite having various forms of existing regression tests from baseline systems, these teams have not yet identified effective methods to adapt them for nondeterministic agent behavior to create test set that cover sufficient runtime scenarios with different nuances.
Verification mechanisms and benefit quantification. Based on the interview case studies, we observe that agent faces two broader evaluation challenges. First, robust verification mechanisms do not always exist. Coding-related agents benefit from strong correctness signals through compilation and test suites, such as software migration agents and site reliability agents in Table 1. However, many agents operate in settings without robust and fast verification. For example, insurance agents receive true signals only through real consequences such as financial losses or delayed patient approvals. These signals arrive slowly and in forms difficult to automate for evaluation. Second, the final benefits of using agents are not always easy to measure. Section 4.1 shows agents target productivity gains because end-to-end time and human hours quantify straightforwardly. Applications with harder-to-measure benefits remain less explored. For example, we observe fewer cases of agents reducing cross-domain knowledge needs or mitigating risks, potentially because benefits manifest indirectly or over longer timeframes.
7.2 Latency Challenges
We examine the degree to which agent execution latency hinders deployment. Survey results indicate that latency represents a manageable friction rather than a hard stop for most teams. Figure 11 shows that only 14.8% of deployed survey agents identify latency as a critical deployment blocker requiring immediate resolution, while the majority (59.3%) report it as a marginal issue, where current latency is suboptimal but sufficient for deployment. We suspect that this tolerance may correlates with the prevalence (15/20 in detailed interview case studies) of asynchronous agent execution paradigm (Section 4.4) and (52.2% from survey) internal user bases (Section 4.3). Notably, we observe a consistent latency distribution across the full survey dataset, including experimental systems (Figure 19(b)). We believe this consistency signals a broader preference for building offline agents, as discussed in Section 4.4.
Interactive agent latency requirements. While latency is not a critical challenge for most agent applications, it remains a critical bottleneck for real-time interactive agents. Two interviewed teams, building voice and specialized chat agents, report continuous engineering efforts to match human conversational speeds. Unlike asynchronous workflows, these systems require seamless turn-taking where delays disrupt the user experience. Achieving fluid real-time responsiveness beyond rigid turn-based exchanges remains an open research question and development challenge.
@@ 7.3 Security and Privacy Challenges
Security and privacy consistently rank as secondary concerns in both of our survey and interviews, with practitioners prioritizing output quality and correctness. Figure 21(a) shows that Compliance and User Trust ranks fourth among challenge categories. Given that Section 4.3 shows 52.2% of systems serve internal employees and many systems with human supervision, this prioritization reflects current deployment environments and requirements rather than dismissing security’s importance.
Data ingestion and handling. Survey results in Figure 12 show that 89.7% of systems ingest information from databases, 65.5% ingest real-time user input, and 51.7% ingest other real-time signals. Notably, 69.0% of systems retrieve confidential or sensitive data, while only 34.5% retrieve persistent public data. Given the high prevalence of sensitive data usage and user inputs, preserving privacy is critical. Our interview case studies reveal that teams address this through legal methods. For example, a team building healthcare agents report relying on standard data-handling practices and strict contractual agreements with model providers to prevent training on their user data.
Security practices. In-depth interview participants describe four approaches to managing security risks through constrained agent design. First, six teams restrict agents to “read-only” operations to prevent state modification. For example, one SRE agent case study generate bug reports and proposes action plans, but leaves the final execution to human engineers. Second, three teams deploy agents in sandboxed or simulated environments to isolate live systems. In one instance, a code migration agent generates and tests changes in a mirrored sandbox, merging code only after software verification. Third, one team builds an abstraction layer between agents and production environments. This team constructs wrapper APIs around production tools, restricting the agent to this intermediate layer and hiding internal function details. Finally, one team attempt to enforce role-based access controls that mirror agent user’s permissions. However, the agent team reports this remains challenging, as agents can bypass these configurations when accessing tools or documents with conflicting fine-grained permissions.
8 Discussion
Building on the quantitative findings in Sections 4–7, we discuss three key trends that characterize the current state of agent engineering. We first examine how practitioners achieve production reliability through architectural constraints (Section 8.1). Next, we highlight how current model capabilities already drive substantial value, revealing additional adoption opportunities (Section 8.2). Finally, we outline concrete open research directions for agent development (Section 8.3).
8.1 Reliability Through Constrained Deployment
A paradox emerges from our data: while nearly 40% of practitioners identify reliability, robustness, and scalability as their primary development concern (Section 7), these systems have successfully reached deployments in production or pilot stages (Section 4). This raises a critical question: how do agents reach production if reliability remains an unsolved challenge? We observe that practitioners ensure reliability via deploying agents with strict constraints on both execution environments and agent autonomy, often combined with close human supervision.
Our in-depth case studies reveals several deployment environment patterns. Some agents operate in read-only mode, never modifying production state directly. For example, SRE agents perform debugging then generate detailed bug reports that engineers review and action. Other agents serve internal users where errors carry lower consequences and human experts remain readily available to correct mistakes (Section 4.3): for example, Slack bot response automation for internal tickets. Systems with write access deploy through sandbox environments where outputs undergo rule-based verification before production integration. Some teams combine read-only access with sandboxes mirroring production environments to further mitigate risk. We observe that these patterns shift the reliability challenge from ensuring correct autonomous agent actions at each step to verifying that final outputs meet acceptable quality thresholds.
Section 5.4 shows 68% of production agents execute fewer than ten steps before requiring human intervention. Organizations deliberately bound agent behavior within specific action spaces through prompting and limited tooling. External-facing systems use particularly restricted agent workflows where customer trust and economic consequences demand tighter control. For example, rather than allowing an agent to autonomously decide if and where to search via tool call, teams employ agentic RAG architectures with pre-determined retrieval steps that restrict the agent to specific document store. We believe this pattern reveals a deliberate engineering trade-off. Production teams balance capability with controllability such that these systems deliver value through well-scoped automation while maintaining reliability guarantees.
An open question remains: how can future agents expand autonomy while maintaining production-level reliability? The field has not established clear pathways for systematically relaxing these constraints while preserving safety guarantees. We observe that one emerging pattern to achieve this goal is the merging of distinct subtasks into a single task. For example, commercial coding assistants adopt test-driven development by merging coding and testing into a unified agent execution flow [ClaudeCode2025, cursor2025]. We believe this pattern of decoupling logical task abstraction from agent execution steps has potential to extend to other domains. The way humans break down tasks into subtasks might not be the right way to organize agent control flow (Figure 22). Progress requires advances in two areas: agent evaluation must move beyond correctness metrics to assess other aspects of reliability (e.g.: five 9s) under varying autonomy levels, and engineering practices that systematically specify and verify bounded operations at different abstraction levels.
8.2 Rich Agent Application Opportunities
Our study reveals an encouraging finding: production agents already operate across more than 26 domains (Figure 2), extending well beyond coding-related agents and chatbots. We find that practitioners extract real value from current model and agent capabilities (Figure 1).
Our detailed case studies show that current frontier models, utilizing prompting strategies alone, already possess sufficient capabilities to cover a diverse range of production use cases. Section 5.1 shows that 70% of deployed survey agents rely on off-the-shelf models without any fine-tuning. This data pattern demonstrates that practitioners enable substantial deployment by leveraging existing model capabilities within well-scoped applications rather than waiting for model improvements. We already see diverse implementations where agents support healthcare workflows, manage business operations, and accelerate scientific discovery in production (Table 1).
Figure 5(a) and our case interviews reveal that relatively little attention has been given to applying agents to software-facing rather than human-facing problems. Most production agents interface directly with users through chat, voice, or interactive workflows. We observe fewer deployments for direct software operation, maintenance, and optimization tasks. We believe opportunities to improve automation in these software-integrated systems remain under-explored, though realizing this potential requires addressing the reliability challenges discussed in Section 8.1.
8.3 Open Research Questions
Beyond the directions we highlight in prior sections, we identify additional research questions based on deployment patterns in RQ1–4.
We show that agents often operate without clean and fast correctness signals in the real-world (Section 7.1). We observe an analogy with silent failures, where catching errors during runtime is challenging and success signals arrive through real consequences like financial loss or customer dissatisfaction. Engineers dedicate substantial effort ensuring agent quality through close human oversight, demonstrating effective practices that enable current deployments. However, the field has yet to reach consensus on effective ways to identify what and when errors or low-quality responses occur. We believe understanding agent failure modes [whydomultiagentllmsystemsfail], developing agent observability tools, and advancing runtime failure mitigation techniques represent important open research directions in addition to benchmarking and CI/CD integration (Section 7.1).
Significant evidence shows interest in post-training for agents [Cursor2025ComposerRL, qu2024recursiveintrospectionteachinglanguage, cao2025skyrlagentefficientrltraining, yuan2025agentrtraininglanguagemodel, Huang2025_biomni, song2024agentbankgeneralizedllmagents, park2025maporlmultiagentpostcotrainingcollaborative, ma2025coevolvingyoufinetuningllm], yet Section 5.1 shows fine-tuning and reinforcement learning remain uncommon in deployment despite potential benefits. Our interviews reveal development effort currently prioritizes making agents work reliably over improving model capabilities for preference optimization, specialization, or cost trade-offs. We speculate this pattern emerges for two reasons: first, supervised fine-tuning and reinforcement learning impose barriers to entry, requiring large amounts of representative data or environments alongside with specialized ML expertise; second, model upgrades introduce complexity when fine-tuned models must adapt to new base versions. We believe the field needs more sample-efficient post-training approaches that remain robust to model upgrades and translate into repeatable engineering practices.
We find in Section 5.4 that 30% of deployed survey agents already adopt inference-time scaling techniques, with the use of multiple models being most popular. From detailed case studies, we find these are commonly simple methods, such as routing distinct subtasks to specialized models. We believe recent advances in generate-verify-search approaches [alphaevolve, barbariansgateaiupending] have significant potential for non-science agent applications as well because they provide stronger reliability guarantees while maintaining controllability and explainability. However, realizing this potential may requires infrastructural support and redesign, such as building simulators and designing verifier. How to adapt agent runtime environments to support such inference-time search technique, and how to effectively utilize them in an online execution settings, are open research questions.
We show in Figure 13 that 93% of current production agents take text or speech in natural language as inputs. Both our survey trends and interviews show high interest in expanding agent modalities to include images, spatiotemporal sequences, video, and scientific data. Interview participants suggest that teams currently focus on applications that are easy to measure and validate. We believe that once these initial deployments work reliably, agent features and capabilities will expand to handle richer input modalities.
9 Conclusion
We present MAP, the first large-scale systematic study characterizing the engineering practices and technical methods behind AI agents actively deployed in production. We analyze data from agent practitioners across 26 domains to answer four research questions regarding the state of real-world agent development:
RQ1: Why build agents? Productivity drives adoption. Organizations deploy agents primarily to automate routine tasks and reduce human task hours, prioritizing measurable efficiency gains over novel capabilities.
RQ2: How are agents built? Simplicity and controllability dominate. Production systems favor closed-source models utilizing manual prompting rather than weight tuning. Architecturally, these systems rely on structured workflows with bounded autonomy and typically execute limited steps before human intervention.
RQ3: How are agents evaluated? Human verification remains the primary method. Practitioners rely heavily on human-in-the-loop evaluation because clean baselines and ground truth datasets are scarce. Automated techniques like LLM-as-a-judge complement human review rather than replace it.
RQ4: What are the challenges? Reliability represents the central development focus. The difficulty of ensuring correctness and evaluating non-deterministic agent outputs drives this friction. Latency and security typically act as manageable constraints rather than hard blockers as engineering workarounds and restricted environments currently manage them.
As “agent engineering” emerges into a new discipline, MAP provides researchers with critical visibility into real-world constraints and opportunities while offering practitioners proven deployment patterns across industries.

### v2

@@ 6.2 Evaluation Methods
Survey data shows human verification dominates evaluation despite automated alternatives. Seventy-four percent rely on human-in-the-loop, where domain experts, operators, or end-users validate the outputs for correctness (Figure 7).
Case study interviews reveal humans verify during both development and production. Development teams work with domain experts to create benchmarks and validate responses (e.g., C15: live systems incident configurations). In production, humans serve as final verifiers—agents suggest actions, experts approve execution (C10: engineers execute on SRE agent recommendation). One team forgoes all automated methods, hand-tuning agent configurations per client through feedback.
Model-based evaluation is second most common (52%), but survey data shows 39% combine it with human verification (Figure 10(b)). All case study teams using LLM-as-a-judge pair them with human review: judges score confidence for each response, routing low-confidence cases to experts while randomly sampling some percentages of high-confidence outputs for ongoing validation (C01, C15).
Survey data shows human-in-the-loop (74%) and LLM-as-a-judge (52%) far exceed rule-based methods (42%). Case study interviews suggest this pattern reflects task complexity—customer support, HR operations, and compliance tasks require nuanced expertise that syntactic checks cannot capture. See §B.3 for detailed method descriptions and co-occurrence patterns.
7 RQ4 Results: What Are The Top Challenges In Deployment?
Survey data shows reliability remains the primary development bottleneck. Thirty-eight percent of practitioners rank “Core Technical Performance”—encompassing reliability, robustness, and scalability—as their top priority in agent development, far exceeding governance (3%) or compliance (17%) (§B.4). We examine three key challenges: how teams achieve reliability, why evaluation remains difficult, and how teams manage security.
7.1 Reliability Challenge
A paradox emerges: if reliability is the primary development bottleneck, how do agent systems reach production? We observe that teams mitigate reliability risks through strict environmental and autonomy constraints.
Interview data shows agents deploy in controlled environments minimizing failure impact: read-only mode where SRE agents generate bug reports for engineer review without modifying production (C16); sandbox verification where systems with write access undergo rule-based checks before production integration (C09, C11-C12); internal deployment serving employees where errors have lower consequences and expert oversight is immediate (§4.3); wrapper APIs restricting agents to abstraction layers that hide production system details (C16); and role-based access controls mirroring user permissions (C15).
Survey data reveals teams deliberately constrain autonomy. Sixty-eight percent execute fewer than 10 steps before human intervention (§5.4). Interviewed teams bound behavior through prompting and limited tooling. External-facing systems use particularly restricted workflows where trust and economic consequences demand control—pre-configured retrieval to specific document stores, mandatory human approval at critical steps, and fixed action sequences.
@@ 7.2 Evaluation Challenges
The patterns observed in §6.2, limited formal benchmarks and dominance of human evaluation (74%), stem from underlying technical challenges revealed in our interviews.
Benchmark scarcity. Interview data reveals three challenges with benchmark creation: (1) Regulated domains lack public data, forcing expensive expert-crafted datasets (C01, C16: months of data collection and labeling). (2) Client-specific customizations make standardized benchmarks infeasible (C04: proprietary toolsets and localized dialogue per deployment), leading teams to default to A/B testing and iterative client feedback.
Real-world tasks are hard to verify. Interview data reveals robust verification mechanisms don’t always exist. Coding agents represent a rare case where verification occurs through compilation and test suites (C09, C12), enabling faster iteration. Most agents operate without fast automated signals. For example, insurance agents receive feedback only through delayed real consequences such as financial losses or patient approval delays (C01). This verification gap may explain both the reliance on human evaluation and the concentration in productivity-focused applications (§4.1), where end-to-end time quantifies straightforwardly while harder-to-measure benefits remain underexplored.
@@ 7.3 Security and Privacy Challenges
Interview data reveals practitioners currently prioritize quality and correctness over security. Interviewed teams share that security is implicitly achieved through systems-level design (e.g., read-only agents) similar to §7.1 approaches. Current deployments also reflect this pattern, for example, 52% serve internal employees (§4.3). Systematic security mechanisms remain an open challenge as deployments expand to higher-stakes external settings.
8 Discussions & Conclusion
We provide the first systematic documentation of technical methods enabling real-world agent deployment. Through 20 in-depth case studies and surveys of 86 deployed systems, we make typically proprietary knowledge accessible to the research community. A central observation emerges from our data: deployment techniques diverge from popular research methods, signaling underexplored research opportunities. We consolidate our 13 findings (F) from RQ1-4 into 3 themes that reveal these research-practice gaps, which we hope these insights can support future agents research.
Surprising deployment patterns. Model brittleness under provider updates drives practitioners away from post-training (F7) despite strong research results. This suggests opportunities in sample-efficient post-training and learning methods that transfer learned behaviors or preserve performance when base models update. Latency-tolerant applications (F4) indicate that inference-time scaling can satisfy agents’ latency requirements, encouraging research in evolutionary search, test-verification architectures, and agent-specific batched inference scheduling. Multi-model coordination (F6) creates opportunities for efficient multi-model inference, orchestration, and routing strategies.
Reliability through system-level design. Practitioners’ preference for simple, controllable methods (F7, F9, F1) over blackbox optimization methods mirrors traditional software engineering practices. Our interviews reveal a distinct agent requirement for production reliability: continuous correctness verification during both development and runtime. This requirement creates research opportunities in observability (runtime monitoring, failure detection), evaluation (correctness metrics, real-world task benchmarks), and failure recovery. ML-focused research opportunities also include training models responsive to system specifications, learning under correctness specifications, and constrained optimization (verifiable actions, planning under bounds).
Human-in-the-loop as design principle. Practitioners deliberately incorporate human-in-the-loop as an architectural component, with agents executing limited steps before human intervention (F3, F9, F12). This currently addresses evaluation limitations (§7.2) and domain knowledge requirements. However, rather than viewing human-in-the-loop as a temporary constraint to overcome, our data suggests it has a potential to endure as a deliberate design principle. While much research emphasizes advancing single-agent autonomy, our findings reveal opportunities in agents augmenting human workers, suggesting research in agent workflows with human oversight, coordination interfaces, and methods for bounded autonomy within human-supervised systems.
As “agent engineering” emerges into a new discipline, MAP provides the research community with empirical insights about deployment realities to collectively advance agents that reliably deliver societal value on real-world tasks.